OpenStudy (anonymous):

Solve the following system using substitution. Write the solution as an ordered pair. 2x + y = 5 8x – y = 45

HanAkoSolo (jamierox4ev3r):

hi there @ami2 to solve with substitution, you'll need to set one equation equal to y for example: 2x+y=5 -2x -2x y= -2x+5 making sense so far?

HanAkoSolo (jamierox4ev3r):

@Ami22 now that we have set the value of y as equal to -2x+5, we can substitute it in to the other equation 8x-y=45 8x-(-2x+5)=45

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Jamierox4ev3r can you answer this for me please 6x - 2y = 50 x + 3/2y = 12